Jessica Wright has always been interested in stocks. She has decided to invest $1,500 once every year into an equity mutual fund that is expected to produce a return of 8 percent a year for the foreseeable future. Jessica is really curious how much money she can reasonably expect her investment to be worth in 22 years. What would you tell her? Round FVA-factor to four decimal places. Round the answer to the nearest cent.
Answer #1

This is the case of Ordinary annuities. $1,500 will invested annually for 22 years. we need to calculate future value of this annuity @ 8% rate of return.

Future value = (Annual investment/rate of return)*[(1+rate of return)no. of years - 1]

Future value = ($1,500/0.08)*[(1+0.08)22 - 1] = $18,750*(1.0822 - 1) = $18,750*(5.4365 - 1) = $18,750*4.4365 = $83,184.38

Jessica can reasonably expect her investment to be worth $83,184.38 in 22 years.